OpenStudy (anonymous):

Find the distance for the following points A ( -2, -3) B (-7, -8). If necessary round your answer to the nearest tenth.

OpenStudy (amistre64):

subtract them to find the distance between your x and y values; then use the pythag thtm to determine distance

OpenStudy (anonymous):

distance = \( \sqrt{ ((-2)-(-7))^2 + ((-3)-(-8))^2 } \) = \( \sqrt{ (2-7)^2 + (3-8)^2} \) = \( \sqrt{5^2 + 5^2} \) = \( \sqrt{50} \) = \( 5 \sqrt{2} \)

OpenStudy (amistre64):

A - B (-2,-3) -(-7,-8) -------- (5 , 5 ) sqrt(5^2 + 5^2) = sqrt(50) = 5sqrt(2)

OpenStudy (anonymous):

5\[what is 5\sqrt{2} \to the nearest tenth\]